OpenStudy (anonymous):

graph the parabola, mark the focus and directrix on graph. y=1/32x2 help please?

OpenStudy (anonymous):

Idk how to graph parabolas

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@bullkidd

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

do you have answer choices?

OpenStudy (anonymous):

no i need to graph it.

OpenStudy (anonymous):

the vertex is (m,k) and p is the distance to both the focus AND the directrix along the axis of symmetry from the vertex (in opposite direction). (focus is in the same direction as the sign of p)

OpenStudy (anonymous):

y= 1/32 x^2 =32y = x^2 =4(8)y = x^2

OpenStudy (anonymous):

so given the vertex is at (0,0) and the axis of symmetry is x=0 then: the focus is (0,8) directrix is y = -8

OpenStudy (anonymous):

what does this look like on a graph? i did it on my calculator and its hard to tell

OpenStudy (anonymous):

thank you